# https://www.informatik.uni-wuerzburg.de/computervision/

### v1

@@ Abstract
Modern smartphone camera quality heavily relies on the image signal processor (ISP) to enhance captured raw images, utilizing carefully designed modules to produce final output images encoded in a standard color space (e.g., sRGB). Neural-based end-to-end learnable ISPs offer promising advancements, potentially replacing traditional ISPs with their ability to adapt without requiring extensive tuning for each new camera model, as is often the case for nearly every module in traditional ISPs. However, the key challenge with the recent learning-based ISPs is the urge to collect large paired datasets for each distinct camera model due to the influence of intrinsic camera characteristics on the formation of input raw images. This paper tackles this challenge by introducing a novel method for unpaired learning of raw-to-raw translation across diverse cameras. Specifically, we propose Rawformer, an unsupervised Transformer-based encoder-decoder method for raw-to-raw translation. It accurately maps raw images captured by a certain camera to the target camera, facilitating the generalization of learnable ISPs to new unseen cameras. Our method demonstrates superior performance on real camera datasets, achieving higher accuracy compared to previous state-of-the-art techniques, and preserving a more robust correlation between the original and translated raw images.
@@ 1 Introduction
  • We propose Rawformer, a fully unsupervised encoder-decoder Transformer-based method designed for raw-to-raw translation facilitating the reuse of neural-based ISPs.
@@
  • We achieve state-of-the-art results on the existing raw-to-raw translation datasets. Additionally, using our method as a pre-processing step for mapping raw images to the camera space used in training neural-based ISPs, we achieve the best results for cross-camera ISP rendering across various datasets.
@@ 2.1 Neural-Based ISP
Several neural-based ISPs methods have been introduced [8, 57, 54, 18], with the majority relying on CNN-based U-Net-like architecture [40]. These include earlier approaches, such as DeepISP [42] and PyNET [21], as well as more recent methods, such as MW-ISPNet [20], AW-Net [12], and LAN [52], which achieve improved image restoration quality by integrating discrete wavelet transform (DWT) and double attention modules (DAM) techniques. Such learnable ISPs are typically trained to map raw images captured by a specific camera model to target high-quality sRGB “ground-truth” images produced by some target camera ISP (usually with higher quality than the mobile native ISP; e.g., a DSLR camera [21, 18]). A significant challenge for neural-based ISPs is the need for paired datasets of raw-sRGB images. Collecting such datasets is tedious and typically involves using specific equipment. Additionally, it may require post-processing techniques to align the raw images with the target sRGB images [21]. The challenge intensifies when a mobile phone manufacturer introduces a new camera model with different characteristics, which happens frequently [47], requiring the recollection of paired raw-sRGB datasets for each new camera model.
@@ 3 Proposed Method
The existing raw-to-raw alternatives map images in a supervised manner (e.g., [35]) or semi-supervised manner (e.g., [4]), which hinders the usability of such methods due to either the need for paired raw-to-raw datasets or limited accuracy. Additionally, for effective reconstruction of the content from the source domain and style from the target domain, domain adaptation methods [37, 59] focus on extracting the local semantics between the domains, but often fall short in comprehending the overall semantics, and context-dependent relationships. In contrast to other methods, our approach in this study handles the raw images in a fully unsupervised manner. We attain this by proposing Rawformer, that effectively leverages the overall semantics via an encoder-decoder architecture in a CycleGAN framework, owing to its ability to maintain intrinsic consistency between the translated images. As illustrated in Fig. 2(a), the generator in Rawformer is specifically designed for efficiently augmenting the content rendering and structural integrity of the generated images via stacked contextual-scale aware downsampler and upsampler blocks. The discriminator (Fig. 3(a)) is modelled for stabilizing the training by blending a cache of the previous crucial discriminator features, thus combating the model collapse and saving redundant computations between different domains. In the following, we provide a comprehensive elucidation of the proposed components.
Figure 2: a) Overview of the generator architecture of Rawformer. The primary components of the generator: b) contextual-scale aware downsampler (CSAD) block, c) condensed query attention block (CQA), d) scale perceptive feed-forward (SPFN) network, e) composite downsampling (CDown) block, f) composite upsampler (CUp) block, and g) contextual-scale aware upsampler (CSAU) block.
3.1 Contextual-Scale Aware Downsampler Block
@@ 3.1.1 Condensed Query Attention Block:
Self-attention has proven its excellence in explicitly modelling the long-range contextual dependencies, where a token is compared with all other tokens enabling seamless information propagation within the network [56]. However, the computational demands of the self-attention pose challenges for generation tasks such as unpaired raw-to-raw translation, where preserving both visual quality and semantic consistency is paramount. To address these concerns and overcome the constraints of computation, we introduce the condensed query attention (CQA) block; see Fig. 2(c). Here, we bring two notable changes in the design of self-attention to improve the overall learning. Firstly, we spatially compress (condense) the spatial dimension of the query vector by applying average pooling followed by a linear projection for confining the crucial information. Following it, we downsample the obtained feature map by factor $r$ in both the dimensions. Thus, instead of directly computing the similarity between the dot product of key-query and value vectors, the condensed queries $(\textbf{Q}_{c})$ act as an intermediate for the similarity comparison as shown in Eq. 1. Secondly, to further strike a balance between the scalability of global and regional range modelling and the computational complexity, the attention is efficiently conducted horizontally and vertically in the network to remove the redundant computations. Formally, the overall operation of CQA block is defined as:
@@ 3.1.4 Style Modulator:
3.2 Contextual-Scale Aware Upsampler Block
@@ 3.3 Cross-Domain Attention-Guided Discriminator
Leveraging batch statistics provides a potent strategy to mitigate mode collapse and foster diversity within the GAN models [41]. The conventional minibatch discrimination techniques for improving the discriminator’s performance necessitate a substantial batch size [24], which could pose challenges due to limitations in GPU hardware. Following the idea in [48], we try to mitigate this issue by decoupling batch size from the minibatch discrimination approach. The core innovation lies in incorporating a cache of the previous discriminator features, functioning as an alternative to larger batch size. It is to be noted that while training, the batch features are stored in 4 different caches $(M)$; real images from the source camera, real images from the target camera, and fake images from both the domains. This concept of the utilization of a memory bank-style cache involves a composite discriminator consisting of a head and a body. The head is designed to capture the batch statistics via a batch normalization layer followed by two 3$\times$3 convolutional layers. It processes a concatenated input (along the batch dimension) comprising the output of the discriminator body for the current minibatch and a historical record of past outputs from the cache, as depicted in Fig. 3(a). The discriminator body comprises of the proposed CQA block followed by strided convolution to check with the highly detailed global features in the distant regions of different domains. This strategic integration of batch statistics, caching mechanism and attention in the discriminator design significantly enhances its capability in handling the multi-resolution image data while stabilising the overall training (Fig. 3 (b) and (c)).
@@ 4.3 Raw-to-Raw Translation Results
Table 1: Raw-to-raw translation results on NUS dataset [9]. The mapping results of Nikon D5200’s raw images to Canon EOS 600D DSLR camera and vice versa using our method and other methods are shown. Here, $\ell_{a}=L_{1}$, $\ell_{b}=L_{1}+VGG$, $\ell_{c}=\mathcal{L}_{pixel-wise}$ (Eq. 3), w, and w/o denote the results with and without training on a corresponding configuration, respectively. Best results are in bold.
@@
CycleGAN [59] 27.32 0.83 0.03 15.40 26.81 0.82 0.03 11.49
UNIT [30] 25.55 0.81 0.03 18.22 24.73 0.79 0.03 18.64
SSRM [4] 32.36 0.93 0.02 6.21 30.81 0.93 0.02 5.95
UVCGANv2 [49] 37.11 0.96 0.02 4.34 37.29 0.96 0.02 4.28
Ours (w/o self-supervised pre-training; $\ell_{a}$ loss) 37.13 0.96 0.02 4.65 37.27 0.96 0.02 4.46
Ours (w/o self-supervised pre-training; $\ell_{b}$ loss) 37.94 0.97 0.02 3.96 37.99 0.96 0.02 3.93
Ours (w/o self-supervised pre-training; $\ell_{c}$ loss) 38.73 0.97 0.02 3.64 38.26 0.97 0.02 3.72
Ours (w self-supervised pre-training; $\ell_{c}$ loss) 41.89 0.98 0.01 2.04 41.37 0.98 0.01 2.53
@@
CycleGAN [59] 24.63 0.71 0.05 14.71 24.37 0.73 0.04 12.88 21.7 19
UNIT [30] 23.91 0.70 0.05 13.23 26.42 0.78 0.04 13.16 32.2 29
U-GAT-IT [26] 28.22 0.89 0.03 5.98 31.04 0.89 0.02 5.04 34.4 35
SSRM [4] 29.65 0.89 0.02 6.32 28.58 0.90 0.02 6.53 20.2 18
UVCGANv2 [49] 36.32 0.94 0.02 4.21 36.46 0.92 0.02 4.73 32.6 31
Ours 40.98 0.97 0.01 2.09 41.48 0.98 0.01 1.99 26.1 26
@@ 8.2 Ablation Study
In the main paper, we presented several ablation studies conducted to validate the decisions made in the proposed design of Rawformer. Here, we present additional ablation experiments performed to further validate the operations of the major components in our Rawformer: condensed query attention (CQA), contextual-scale aware upsampler (CSAU) block and the contextual-scale aware downsampler (CSAD) blocks. Tables 9, and 10 demonstrates the impact of incorporating the spatial compression (where and how), on the attention vectors of CQA block. The results of the ablation studies shown in Tables 11, and 12 proves the efficacy of our designed composite upsamplers and downsamplers. All these ablation studies clearly reveal, that the proposed design with the inclusion of all the components achieves the best results across all quantitative metrics.

### v2

@@ Abstract
Modern smartphone camera quality heavily relies on the image signal processor (ISP) to enhance captured raw images, utilizing carefully designed modules to produce final output images encoded in a standard color space (e.g., sRGB). Neural-based end-to-end learnable ISPs offer promising advancements, potentially replacing traditional ISPs with their ability to adapt without requiring extensive tuning for each new camera model, as is often the case for nearly every module in traditional ISPs. However, the key challenge with the recent learning-based ISPs is the urge to collect large paired datasets for each distinct camera model due to the influence of intrinsic camera characteristics on the formation of input raw images. This paper tackles this challenge by introducing a novel method for unpaired learning of raw-to-raw translation across diverse cameras. Specifically, we propose Rawformer, an unsupervised Transformer-based encoder-decoder method for raw-to-raw translation. It accurately maps raw images captured by a certain camera to the target camera, facilitating the generalization of learnable ISPs to new unseen cameras. Our method demonstrates superior performance on real camera datasets, achieving higher accuracy compared to previous state-of-the-art techniques, and preserving a more robust correlation between the original and translated raw images. The codes and the pretrained models are available at https://github.com/gosha20777/rawformer.
@@ 1 Introduction
  • We propose Rawformer, a fully unsupervised encoder-decoder Transformer-based method designed for raw-to-raw translation facilitating the reuse of neural-based ISPs. Rawformer achieves state-of-the-art results in raw-to-raw translation, leading to better generalization for cross-camera learnable ISP rendering.
@@ 2.1 Neural-Based ISP
Several neural-based ISPs methods have been introduced [8, 62, 58, 21, 63, 26, 17], with the majority relying on CNN-based U-Net-like architecture [44]. These include earlier approaches, such as DeepISP [46] and PyNET [24], as well as more recent methods, such as MW-ISPNet [23], AW-Net [12], and LAN [56], which achieve improved image restoration quality by integrating discrete wavelet transform (DWT) and double attention modules (DAM) techniques. Such learnable ISPs are typically trained to map raw images captured by a specific camera model to target high-quality sRGB “ground-truth” images produced by some target camera ISP (usually with higher quality than the mobile native ISP; e.g., a DSLR camera [24, 21]). A significant challenge for neural-based ISPs is the need for paired datasets of raw-sRGB images. Collecting such datasets is tedious and typically involves using specific equipment. Additionally, it may require post-processing techniques to align the raw images with the target sRGB images [24]. The challenge intensifies when a mobile phone manufacturer introduces a new camera model with different characteristics, which happens frequently [51], requiring the recollection of paired raw-sRGB datasets for each new camera model.
@@ 3 Proposed Method
Existing raw-to-raw methodologies either rely on supervised approaches requiring paired datasets e.g., [39]) or semi-supervised approaches (e.g., [4]) with limited accuracy. Additionally, domain adaptation techniques (e.g., [41, 65]), which aim to extract local semantics between source and target domains, often fail to capture overall semantics and context-dependent relationships. Our proposed method, Rawformer, addresses these limitations by processing raw images in a fully unsupervised manner. Utilizing an encoder-decoder architecture within a CycleGAN framework, Rawformer effectively captures global semantics and maintains intrinsic consistency between translated images. The generator (Fig. 2(a)) in Rawformer employs contextual-scale aware downsampler and upsampler blocks to enhance content rendering and structural integrity. The discriminator (Fig. 3(a)) stabilizes training by incorporating key discriminator features, thereby preventing model collapse and reducing redundant computations. In the following, we provide a comprehensive elucidation of the proposed components.
Figure 2: a) Overview of the generator architecture (a) of Rawformer. The primary components of the generator: b) contextual-scale aware downsampler block (CSAD), c) condensed query attention block (CQA), d) scale perceptive feed-forward network (SPFN), e) composite downsampling block (CDown), f) composite upsampler (CUp) block, and g) contextual-scale aware upsampler block (CSAU).
3.1 Generator’s Encoder Network
@@ 3.1.1 Condensed Query Attention Block:
Self-attention is highly effective in modeling long-range contextual dependencies by comparing each token with all others, facilitating seamless information propagation within the network [61]. However, its computational demands pose challenges for tasks like unpaired raw-to-raw translation, where maintaining both visual quality and semantic consistency is crucial. To address these concerns and overcome the constraints of computation, we introduce the condensed query attention (CQA) block; see Fig. 2(c). Here, we make two key modifications to self-attention to enhance learning efficiency. Firstly, we spatially compress the query vector using average pooling followed by linear projection, and downsample the resulting feature map by a factor of $r$ to confine the crucial information. This results in condensed queries,$(\textbf{Q}_{c})$ used for similarity comparison, as shown in Eq. 1. Secondly, to balance global and regional modeling with computational efficiency, attention is applied horizontally and vertically to eliminate redundant computations. The CQA block operation is defined as:
@@ 3.1.4 Style Modulator:
3.2 Generator’s Decoder Network
@@ 3.3 Cross-Domain Attention-Guided Discriminator
Leveraging batch statistics is an effective strategy to mitigate mode collapse and enhance diversity in GAN models [45]. Traditional minibatch discrimination techniques require large batch sizes [28], posing challenges due to GPU limitations. Inspired by [52], we address this by decoupling batch size from the minibatch discrimination approach. Our core innovation involves using a cache of previous discriminator features as an alternative to larger batch sizes. During training, batch features are stored in four separate caches $(M)$; real images from the source and target cameras, and fake images from both domains. This memory bank-style cache utilizes a composite discriminator with a head and body. The head captures batch statistics via batch normalization and two 3$\times$3 convolutional layers, processing a concatenated input of current and historical discriminator outputs from the cache, as shown in Fig. 3(a). The discriminator body includes a CQA block followed by strided convolution to handle detailed global features across domains. This integration of batch statistics, caching, and attention mechanisms significantly enhances the discriminator’s ability to manage multi-resolution image data and stabilize the overall training (Figs. 3 b and c).
3.4 Self-Supervised Pre-Training
Self-supervised pre-training optimizes the initial weights of a network by engaging it in auxiliary tasks formulated from the inherent structure of the original data. In general, despite modifications in the cross-domain attention-guided discriminator, GAN training still suffers from mode collapse, where the generator repeatedly tends to produce a limited set of images. Thus, we opt to pre-train the generator of our Rawformer using self-supervised inpainting tasks to enhance initial weight optimization and mitigate this mode collapse problem. More details about the mode of pretraining is provided in the supplementary (Section 1) of the paper.
@@ 4.3 Raw-to-Raw Translation Results
Table 1: Raw-to-raw translation results on NUS dataset [9]. The mapping results of Nikon D5200’s raw images to Canon EOS 600D DSLR camera and vice versa using our method and other methods are shown. Here, $\ell_{a}=L_{1}$, $\ell_{b}=L_{1}+VGG$, $\ell_{c}=\mathcal{L}_{pixel-wise}$ (Eq. 3), w, and w/o denote the results with and without training on a corresponding configuration, respectively. Additionally, we show the GFLOPs of each method for the 256×256 input image. Best results are in bold.
@@
CycleGAN [65] 27.32 0.83 3.21 15.40 26.81 0.82 3.33 11.49 36.8
UNIT [34] 25.55 0.81 3.17 18.22 24.73 0.79 3.23 18.64 47.3
U-GAT-IT [30] 29.41 0.87 3.02 7.12 27.86 0.85 3.29 8.01 62.9
SSRM [4] 32.36 0.93 2.41 6.21 30.81 0.93 2.47 5.95 37.1
UVCGANv2 [53] 37.11 0.96 2.32 4.34 37.29 0.96 2.28 4.28 69.6
Ours (w/o self-supervised pre-training; $\ell_{a}$ loss) 37.13 0.96 2.29 4.65 37.27 0.96 2.31 4.46 52.7
Ours (w/o self-supervised pre-training; $\ell_{b}$ loss) 37.94 0.97 2.21 3.96 37.99 0.96 2.23 3.93 52.7
Ours (w/o self-supervised pre-training; $\ell_{c}$ loss) 38.73 0.97 2.19 3.64 38.26 0.97 2.22 3.72 52.7
Ours (w self-supervised pre-training; $\ell_{c}$ loss) 41.89 0.98 1.46 2.04 41.37 0.98 1.41 2.53 52.7
@@
CycleGAN [65] 24.63 0.71 5.40 14.71 24.37 0.73 4.34 12.88 21.7 19
UNIT [34] 23.91 0.70 5.15 13.23 26.42 0.78 4.27 13.16 32.2 29
U-GAT-IT [30] 28.22 0.89 3.11 5.98 31.04 0.89 2.39 5.04 34.4 35
SSRM [4] 29.65 0.89 2.22 6.32 28.58 0.90 2.43 6.53 20.2 18
UVCGANv2 [53] 36.32 0.94 2.09 4.21 36.46 0.92 2.02 4.73 32.6 31
Ours 40.98 0.97 1.33 2.09 41.48 0.98 1.21 1.99 26.1 26
@@ 5 Conclusion and Future Work
Acknowledgments
This work was partly supported by The Alexander von Humboldt Foundation.
@@ 8.2 Ablation Study
In the main paper, we presented several ablation studies conducted to validate the decisions made in the proposed design of Rawformer. Here, we present additional ablation experiments performed to further validate the operations of the major components in our Rawformer: condensed query attention (CQA), contextual-scale aware upsampler (CSAU) block and the contextual-scale aware downsampler (CSAD) blocks. Tables 9, and 10 demonstrates the impact of incorporating the spatial compression (where and how), on the attention vectors of CQA block. The results of the ablation studies shown in Tables 11, and 12 proves the efficacy of our designed composite upsamplers and downsamplers. Additionally, to prove that the different hierarchical levels in our proposed SPFN module offer a more intuitive approach for the impact of capturing local image structure across different scales, we provided an ablation study as shown in Table 13. All these ablation studies clearly reveal, that the proposed design with the inclusion of all the components achieves the best results across all quantitative metrics.
@@
Table 13: Ablation results of different modifications of SPFN.
Task PSNR SSIM MAE (×$\times 1e-2$) $\Delta$E
Multi-scale SPFN 41.89 0.98 1.26 2.04
Single-scale SPFN 41.11 0.97 1.70 2.57
9 Comparison with SoTA approaches
Our method represents the first practical and precise solution to this challenge, relying solely on unpaired sets of raw images. While our network design draws inspiration from existing techniques like channel dependencies from SENet [20] and depthwise convolutions from MobileNet [19], our goal is to develop an efficient network capable of achieving precise unpaired image translation with minimal resource requirements. Given the typical deployment of such tasks on resource-constrained devices, our modifications aim to not only reduce resource demands compared to alternatives mentioned above, but also to enhance accuracy. As demonstrated in the main paper (Tables 1-3), our design outperforms several alternative methods (e.g., UVCGANv2). Even when substituting certain blocks in our design with alternative ones, our proposed design achieves better results with reduced computations. For instance, our discriminator utilizes an attention mechanism with caching, yielding improved results compared to the UVCGANv2 discriminator: 41.89 dB vs. 40.51 dB on the Canon-to-Nikon set (CNS). Our efficient CQA block, which performs spatial compression/condensation operations on query vectors, achieves better results compared to traditional self-attention: 41.89 dB vs. 39.93 dB on CNS with a 50.7% reduction in FLOPs (also see Table 2 in supp. materials). Using our single resolution SPFN block leads to improved results compared to mutli-resolution HRNet [60] feedforward block : 41.89 dB vs. 41.77 dB on CNS with $\sim$87reduction in FLOPs. We will further refine our main contributions and clarify our design motivation, as suggested by R1& R4, respectively thus eventually highlighting the ability of accurate raw-to-raw with unpaired data.
@@ 10 Additional Results
Additionally, we conducted experiments to test the impact of our proposed method on rendering images using Adobe Lightroom. For this experiment, we used the Raw-to-Raw dataset [4]. We compared raw images captured by a Samsung smartphone, rendered by Adobe Lightroom with metadata stored in raw DNG images captured by an iPhone, against Samsung raw images that were mapped to the iPhone raw space using our Rawformer. We performed a similar comparison on rendering iPhone images with Samsung DNG metadata, both with and without our mapping.
Table 15 presents the results, along with those of Adobe Lightroom rendering raw/metadata taken from the same phone, which represents the best case where no mapping is needed. The results are compared against the ground-truth sRGB images of the target camera. As shown in Table 15, we achieve consistent results similar to those in the referenced paper. The empirical results demonstrate that besides achieving superior performance on learnable image signal processors (ISPs), Rawformer also exhibits a competitive edge when evaluated against commercial ISPs, like Adobe Lightroom. This indicates its robust adaptability and effectiveness across diverse ISP implementations, highlighting its potential for both research and commercial applications.
Table 15: Results (PSNR/SSIM) via Lightroom (LR). Off-diagonal results: mapping source camera raw to target camera using Rawformer before rendering with LR using target camera’s metadata.
Target camera
Source camera Samsung iPhone
Samsung 22.76/0.80 21.98/0.77
iPhone 21.42/0.76 22.08/0.79